### v2

# Out of the Box, into the Clinic? Evaluating State-of-the-Art ASR for Clinical Applications for Older Adults

Abstract
Voice-controlled interfaces can support older adults in clinical contexts – with chatbots being a prime example – but reliable Automatic Speech Recognition (ASR) for underrepresented groups remains a bottleneck. This study evaluates state-of-the-art ASR models on language use of older Dutch adults, who interacted with the Welzijn.AI chatbot designed for geriatric contexts. We benchmark generic multilingual ASR models, and models fine-tuned for Dutch spoken by older adults, while also considering processing speed. Our results show that generic multilingual models outperform fine-tuned models, which suggests recent ASR models can generalise well out of the box to realistic datasets. Furthermore, our results suggest that truncating existing architectures is helpful in balancing the accuracy-speed trade-off, though we also identify some cases with high WER due to hallucinations.
Out of the Box, into the Clinic? Evaluating State-of-the-Art ASR for Clinical Applications for Older Adults
Bram van Dijk1, Tiberon Kuiper1, Sirin Aoulad si Ahmed1, Armel Levebvre1, Jake Johnson1, Jan Duin1, Simon Mooijaart1, Marco Spruit1,2 1Leiden University Medical Centre 2Leiden Institute of Advanced Computer Science b.m.a.van_dijk@lumc.nl
1 Introduction
Although there is a surge of interest in AI-driven applications like chatbots in the health domain (Guo et al., 2024; Huo et al., 2025), tailoring them to groups underrepresented in AI research remains a challenge. Older populations are one example: because they often have different needs and preferences when interacting with AI (van Dijk et al., 2025; Klaassen et al., 2025), their involvement in system development is key for building clinically relevant systems in geriatrics, the field in healthcare concerned with the health of older adults. This population is increasing in size globally (World Health Organisation, 2023), while personnel shortages in healthcare become pressing (European Union, 2023), yet addressing these challenges with AI warrants systems that align with older populations.
Voice control is a key element in chatbots in geriatrics, as older individuals may struggle with small fonts, icons, and typing text in standard interfaces (Khamaj, 2025). Yet, implementing voice control is not obvious, as the performance of Automatic Speech Transcription (ASR) systems depends on the representation of older people in training data, and on their articulation, speech volume and experience with technology (Klaassen et al., 2025). Still, evaluation of state-of-the-art ASR systems on realistic data of older individuals is lagging behind.
In this short paper we evaluate recent ASR models on older individuals’ language use in interaction with Welzijn.AI. This is a new digital platform for older users, that currently features a prototype chatbot to converse about clinically relevant topics like quality of life and frailty. Audio data of 10 Dutch older individuals interacting with Welzijn.AI was transcribed using generic multilingual ASR models (Whisper and Voxtral), and models fine-tuned specifically for Dutch or older Dutch populations (Whisper and Wav2Vec). Our findings show that i) generic multilingual models outperform fine-tuned models, and also suggest that ii) truncating larger architectures may help striking a good balance between accuracy and speed. These findings also hold for a similar subset of the Common Voice (CV) dataset we use as benchmark.
Figure 1: Interface of the chatbot with an example conversation. Users press the purple button to respond, and their speech is transcribed and rendered on the screen. Chatbot responses were read out using a text-to-speech model. The ‘Appearance’ button is for adjusting looks, e.g. font size, ‘Scores’ for consulting information extracted on quality of life and frailty, and ‘Settings’ for choosing different ASR models. In this paper we focus only on data resulting from interaction with this prototype.
2 Background
Though work on evaluating ASR models for older persons exists, extrapolating findings to realistic contexts is hard. Performance of ASR models sometimes remains implicit in downstream use, for example through evaluation of overlap of linguistic features extracted from model vs. human-generated transcripts (Naffah et al., 2025), or by using only part of ASR models (like the audio encoder) in predicting cognitive impairment in older persons (Agbavor and Liang, 2024). Other work tailors ASR systems to older individuals by drawing on additional databases of individual users Xu et al. (2025), which is not always feasible in a clinical context due to privacy concerns. However, Xu et al. (2025) also show that fine-tuning generic multilingual models with speech from older persons increases performance. The work by Shekoufandeh et al. (2025) explores this further by fine-tuning Whisper as common standard ASR model, (Radford et al., 2023) on the Dutch JASMIN-CGN dataset. This dataset includes language use of older users in a human-machine interaction setting designed to evoke speech phenomena associated with ASR errors, like additional emphasis, hesitation and frustration Cucchiarini and Van hamme (2012), which are also relevant to Welzijn.AI.
3 Materials and methods
Ten older Dutch volunteers ($\geq$ 65 years) were recruited from the geriatrics outpatient clinic of the Leiden University Medical Center. The geriatric department’s ethics review board approved this study. Participants interacted individually with the chatbot, with one experimenter standing by for help.
An impression of the chatbot is given in Figure 1. Here we discuss only the aspects relevant for supporting conversations, and refer for details on architectures that inspired this system to van Dijk et al. (2025). The chatbot was driven by the Llama3.3 70B Instruct language model Dubey et al. (2024), prompted to have a conversation structured around the EQ-5D and Clinical Frailty Scale (Brooks, 1996; Rockwood et al., 2005). These are common geriatric instruments to assess health-related quality of life and level of frailty. The instruments retrieve information about mobility, mental wellbeing, pain, physical independence, and so on. They can be used in questionnaire form, or in natural conversations with clinicians, or chatbots (topics emerge in the example conversation in Figure 1).
The chatbot ran on a laptop, and interactions took 5-10 minutes and were recorded with a phone, anticipating using the app on handheld devices and relying on built-in microphones. The ASR model used in the chatbot was Whisper large V3, which in early testing we found to work best.
Type Example
Orthographic Hi, I am uh… feeling great today.
Orthographic_clean Hi I am feeling great today
Normalised hi i am feeling great today
Table 1: Examples of gold transcript types.
HuggingFace ID Params
mistralai/Voxtral-Mini-3B-2507 4.68B
golesheed/whisper-native-elderly-9-dutch 1.54B
openai/whisper-large-v2 1.55B
openai/whisper-large-v3 1.55B
golesheed/wav2vec2-xls-r-1b-dutch-3 963M
openai/whisper-large-v3-turbo 809M
openai/whisper-medium 769M
openai/whisper-small 244M
Table 2: Models used for our ASR task.
HuggingFace ID Welzijn.AI chatbot Mozilla Common Voice
Orth. Orth_c Norm. Time Orth. Orth_c Norm. Time
mistralai/Voxtral-Mini-3B-2507 .17 .11 .09 3.75 .05 .04 .04 3.72
golesheed/whisper-native-elderly-9-dutch .40 .22 .14 4.59 .29 .18 .07 4.07
openai/whisper-large-v2 .19 .12 .10 4.69 .05 .04 .04 4.24
openai/whisper-large-v3 .12 .07 .06 3.41 .04 .03 .03 3.71
golesheed/wav2vec2-xls-r-1b-dutch-3 .49 .37 .36 .99 .30 .19 .19 .89
openai/whisper-large-v3-turbo .16 .10 .08 1.43 .06 .04 .04 1.47
openai/whisper-medium .19 .13 .11 2.35 .07 .06 .06 2.48
openai/whisper-small .26 .18 .17 1.11 .12 .10 .10 1.19
Table 3: WER (avg. number of errors per reference word) for Orthographic (Orth.), Orthographic_clean (Orth_c), and Normalised (Norm.) gold transcripts. Processing time (Time) in avg. seconds per input. Best results in bold.
Recorded user speech was after the interactions segmented and separated from chatbot responses using PyDub [^1] https://github.com/jiaaro/pydub and pyannote (Bredin, 2023). We obtained 199 segments with an average length of 3̃ seconds, due to the turn-taking nature of the conversation. Our sample totals 11 min. and 15 sec., which is small, but still a valuable dataset as realistic clinical data is not easy to obtain. Besides chatbot data, we also drew 200 random samples of the Common Voice (CV) dataset (Ardila et al., 2020), of Dutch older individuals ($\geq$ 60), which totals 17 min., and 37 sec. This concerns read speech hence is an easier benchmark for ASR models.
For obtaining gold standard (human) transcriptions, segments were transcribed with the default Whisper large V3 model, and subsequently manually corrected by the first author. Since the Word Error Rate (WER) as standard metric in the ASR field is sensitive to fillers, capitalisation, and punctuation, and since the ‘right’ gold transcription depends on the use case, we created three types of human and model transcriptions as visible in Table 1: orthographic (including punctuation, fillers, and capitalisation) orthographic_clean (only capitalisation, as also used in Figure 1), and normalised (no capitals, fillers, or punctuation). Orthographic transcription is useful in that it provides additional structure, though speech content is often arguably sufficiently preserved in normalised transcriptions, with orthographic_clean transcription striking a balance between the strictest and most flexible evaluation scenarios.
Table 2 shows the ASR models included in our experiments. We focus on the Whisper model family as the current standard in the field, and include the Voxtral 4.68B model (Liu et al., 2025) as fitting contender regarding size. We include the Whisper model fine-tuned for Dutch older individuals by Shekoufandeh et al. (2025), and also included a work-in-progress Wav2Vec model fine-tuned for Dutch in general (Baevski et al., 2020) by the same authors, as older but lightweight and fast contender. We note that these fine-tuned models do not output capitals, punctuation, or fillers by default, hence evaluate them with normalised transcripts. We opt for these models as they are all small (compared to Voxtral 24B for example) and open weights, hence suitable for downstream applications. We take processing time into account since a good balance between accuracy and speed is key in many applications, so include models of different sizes.
Our code for the ASR pipeline is available. [^2] https://anonymous.4open.science/r/asr_pipelines-361F All experiments were carried out on a Macbook Pro M1 16GB using the HuggingFace ecosystem for ASR and PyTorch’s MPS GPU acceleration backend. Due to privacy restrictions we cannot share recordings nor transcripts of the interactions.
# Model Prediction Reference WER
1 openai/whisper-large-v2 hartelijk bedankt voor het kijken en tot de volgende keer gaat wel 5
thank you cordially for watching and until next time it’s okay
2 golesheed/whisper-native-elderly-9-dutch poet hem in oranje goedemiddag 4
loot him in orange good afternoon
3 openai/whisper-large-v2 ik denk dat het weer zo is als het altijd is ongeveer zoals altijd 3.33
I think that it is as it always is roughly as usual
4 openai/whisper-medium ik ben benieuwd goedemiddag 3
I am curious good afternoon
5 openai/whisper-medium voor de wereld goedemiddag 3
for the world good afternoon
6 mistralai/Voxtral-Mini-3B-2507 groen de medaillon goedemiddag 3
green the medallion good afternoon
7 golesheed/wav2vec2-xls-r-1b-dutch-3 goede midda goedemiddag 2
good midda good midday
Table 4: Example inputs with high WER. English (literal) translations in italics.
4 Results
WER and processing time per model are given in Table 3. Regarding WER, on both the Welzijn.AI and CV data, Whisper V3 large as generic multilingual model outperforms all others, even for normalised transcripts as most flexible evaluation scenario. In terms of processing time, on both datasets, the Wav2Vec model fine-tuned on Dutch data is the fastest, but also the least accurate.
For chatbot applications like Welzijn.AI, understanding the trade-off between WER and processing time is crucial, and we visualise performance on normalised transcripts in Figure 2. Here we see that the Whisper V3 large turbo model, which is essentially Whisper V3 large but with a truncated decoder 1/8 its size, strikes the best balance. The ‘nearest’ improvement in WER concerns Whisper V3 large, which is almost three times slower, while the ‘nearest’ improvement in processing time comes from Whisper small, at the expense of more than twice its WER. Figure 2, also shows CV is an easier dataset regarding WER, though not all models process these data faster. For CV the same observations hold regarding the ‘nearest’ improvements in WER and processing times compared to Whisper V3 large turbo: they lead to considerable drops in speed and accuracy.
To get an idea of problematic inputs for ASR models, we give examples of inputs with the highest WER in Table 4, which as expected come only from interaction data. Some seem entirely hallucinated (1, 4 and 5), some seem to align slightly better with phonemes (2, 6), or original meaning (3). High WER occurs for all models, which aligns with earlier documented unpredictability in state-of-the-art ASR models (Koenecke et al., 2024), that should inform their development and deployment in high-stakes contexts.
Figure 2: Overview of accuracy vs. processing time.
5 Discussion and conclusion
We evaluated state-of-the-art ASR models on transcribing language use of older individuals interacting with the Welzijn.AI chatbot, which was designed for geriatrics. We included various generic multilingual models as well as models fine-tuned on language use of older Dutch populations and general Dutch. We found that on both chatbot interaction and CV data, generic multilingual models perform better than fine-tuned models, with Whisper V3 large emerging as best model. Interestingly, its truncated variant Whisper V3 large turbo struck the best balance between accuracy and processing speed. This is useful from a systems development perspective, since these models may perform well out-of-the-box, without the need for training architectures with smaller numbers of parameters for faster processing from scratch, or the need for additional data for fine-tuning to a specific group or task. Still, inspection of transcription errors across the board revealed some challenging inputs with high WER due to what seem hallucinations, where there is little connection between the reference qua semantics or phonemes and transcript.
Other evaluations of state-of-the-art ASR models on the entire Dutch subset in CV data reported WERs of .06 by Voxtral 4.68B (Liu et al., 2025), and .04 by Whisper V3 large. [^3] Results of latest model reported on https://github.com/openai/whisper, August 6 2025.Assuming the strictest evaluation scenario (orthographic), our results are similar: .04 for Whisper V3 large and .06 for Voxtral 4.68B. Our Welzijn.AI data is rare, hence WER results are harder to put in perspective; still, though the WER on orthographic transcripts for Welzijn.AI data triples for the same Voxtral and Whisper models, from the perspective of the difference between read speech and interaction data, this is not a dramatic loss of performance.
6 Limitations
Though dedicated GPUs in HPC clusters have higher bandwidths and are faster, hence the default choice in experiments like ours, unified architectures (such as provided by Apple’s Silicon M series) are receiving more attention nowadays due to their benefits in terms of latency, energy consumption and system footprint, and are recognized as efficient and competitive alternatives to dedicated GPUs (Hübner et al., 2025; Kenyon and Capano, 2022). So faster processing times could probably be attained by using dedicated GPUs or by using a more recent M-chip. Still, we anticipate a scenario where one device hosts multiple smaller AI models to do different tasks, for which our setup can provide a good lower bound.
Also, though we tried to make comparison fair for different ASR models by evaluating with different kinds of transcripts, developing further transcript normalisations to take ‘acceptable errors’ into account, e.g. writing numbers in digits or letters (‘8’, ‘eight’), were beyond the scope of the current work. This means that there can be some noise in our performance estimates.